# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

### v1

@@ 7 Light Curve modeling
To estimate the explosion parameters and infer the properties of CSM around SN 2023ixf, we performed numerical light curve modeling of SN 2023ixf. We used one-dimensional multi-frequency radiation hydrodynamics code STELLA (Blinnikov et al., 1998, 2000, 2006) for this purpose. Moriya et al. (2023) computed light-curve models of SNe II with various progenitor masses, explosion energies, and CSM density profile and its extent using STELLA. The model grid adopts the solar-metallicity RSG progenitor models with the ZAMS masses of 10, 12, 14, 16, 18 from Sukhbold et al. (2016) and covers explosion energies from to . The CSM is constructed by assuming the mass-loss rates of with a terminal wind velocity of . The density structure just above the progenitors is characterized by the parameter ranging from to . We refer the reader to Moriya et al. (2023) for additional details regarding the model grid. The extensive details regarding the model comparison will be presented in Moriya & Singh (2024).
@@
Using the pre-computed model grid, we first looked for the best matching models in the g and r bands by searching for models with the minimum . We found that the models from a low-mass progenitor () best fit the light curves of SN 2023ixf. The best fitting models have the explosion energy of around , the mass of around , the mass-loss rate of around 10-2 , the CSM radius of around , and of around . We performed extra numerical light curve calculations and found that the model with the CSM radius of matches the observed dataset, including the photospheric velocity (shown in dashed lines in Fig. 13). The mass of the confined CSM below in this model is . This extra model assumes that is uniformly mixed in the entire ejecta. We found that the fully mixed model better fits the transition from the plateau phase to the tail phase. However, they also lead to an excess emission in the redder bands during the late plateau phase as starts diffusing out earlier than in the case of a centrally concentrated 56Ni. A much more prominent effect of -mixing was seen in the case of SN 2009ib (Takáts et al., 2015) and SN 2016gfy (Singh et al., 2019a). Since the 56Ni synthesized in SN 2023ixf is high in comparison to normal SNe II (0.03 , Anderson et al., 2014a), the late-plateau luminosity bump is much more prominent in the redder bands of our models.
@@
We observed a notable decline in the UV flux around 40 days in the best-fitting model (having only an inner dense CSM) presented above in Fig.13. To address this decrease and account for the sustained UV excess in the late phase of SN 2023ixf, we modeled the light curves with an extended CSM component situated right outside the inner CSM from to . This additional CSM component possesses a mass of 0.025 and has a density structure proportional to . The low-density extended CSM component helps sustain a prolonged UV brightness until the early nebular phase. Assuming a constant wind velocity of , the outer CSM component is shaped by an average mass-loss rate of 10-4 . The mass-loss rate should increase towards the time of the explosion to form a density structure proportional to . We found that the shallower CSM structure, such as that proportional to , gives a UV flux excess much higher than observed.
@@
SN 2023ixf has a higher peak luminosity, higher 56Ni-mass, and a higher photospheric velocity than a typical Type II SN. The peak -band luminosity of SN 2023ixf is -18.2 mag, which is 4 times brighter than compared to the luminosity of a typical Type II SN, i.e., -16.7 mag (Anderson et al., 2014a). The 56Ni-mass of SN 2023ixf is 0.059 0.001 which is 80% higher than the 56Ni-mass estimate of a typical Type II SN, i.e., 0.033 from (Anderson, 2019). Previous studies on SNe II (Hamuy, 2003; Pejcha & Prieto, 2015; Gutiérrez et al., 2017b) have shown that more energetic explosions lead to higher photospheric velocities and a higher 56Ni-yield in the SN. The estimated 56Ni-mass does indicate that SN 2023ixf is a highly energetic event and is in concordance with the high explosion energy of 2 10 estimated from the light curve modeling. Other works on numerical modeling of SN 2023ixf also suggest high explosion energies. Bersten et al. (2024) estimate 1.2 foe as explosion energy for a 12 RSG with 10.9 as the final progenitor mass. Hiramatsu et al. (2023) estimated 1 foe as explosion energy while exploring 2 foe as well for a 12 progenitor with 11 as the final progenitor mass.
@@
In our case, where we have the best match with a lower mass progenitor, slightly higher explosion energies must be required to match the observed light curves. Although we fit both photospheric velocities and light curves simultaneously, our mass and radius of the progenitors are fixed since we adopt progenitor models from Sukhbold et al. (2016). We cannot lift degeneracies between the ejecta mass, radius, and explosion energy (Goldberg et al., 2019). However, the high-explosion energy cannot entirely explain the peak luminosity of SN 2023ixf. It is enhanced further by early interaction with CSM. The photospheric velocity of SN 2023ixf is 20% faster than a proto-typical Type II SN at 50 d (see Section 5.4), and a power-law fit to photospheric velocity evolution returned exponent -0.47 0.04, which is slower than the average value derived from a large sample of SNe IIP (–0.581 0.034) in Faran et al. (2014). However, since SN 2023ixf is a short-plateau SN, the slightly low-ejecta mass could be one reason for its higher photospheric velocity (Teja et al., 2022).
@@ 8.1 Timeline of significant epochs for all historic observations
Epoch of Polarization Peak A | 1.4 d | 4, This paper | |
@@
Peak in UV Luminosity (blue peak) | 4.5 0.5 d | This paper | |
@@
Emergence of Broad absorption | 7 d | This paper | |
End of Flash-Ionisation Phase | 8 d | 1, 5, This paper | |
@@
Epoch of Radio Detection (10 GHz) | 29.2 d | 11 |
(1) Teja et al. (2023a) (2) Smith et al. (2023); (3) Zimmerman et al. (2023); (4) Vasylyev et al. (2023); (5) Jacobson-Galan et al. (2023); (6) Grefenstette et al. (2023); (7) Itagaki (2023); (8) Bostroem et al. (2023); (9) Chandra et al. (2023); (10) Panjkov et al. (2023); (11) Matthews et al. (2023)
In its infancy, SN 2023ixf showed a rapid evolution spearheaded by the appearance of an increase in ionization in the early flash-ionization phase, accompanied by a rapid ascent in the early UV flux. Additionally, the increase in the color temperature to approximately 35,000 K until around 2.2 days signifies that the shock breakout occurred within a confined, dense CSM. The epoch of peak UV luminosity (and the bolometric peak) at 4.5 d is synonymous with the peak of the He II line flux of Zimmerman et al. (2023), which traces the strength/flux of the flash-ionization features. The change in the line profile of the narrow feature and the drop in its strength seen in the high-resolution spectroscopy of SN 2023ixf (Smith et al., 2023) also happens around 4.4 d, and is synonymous with UV peak. This indicates that the thermal heating and ionization continued beyond the shock breakout (2.2 d). The emergence of a CDS (Chugai, 2009) within the post-shock CSM and the decelerated SN ejecta is probably contributing to the photo-ionization and prolonged heating observed in the flash-ionized features of SN 2023ixf. During the breakout phase, the radius of thermal emission remained relatively constant at (2.0 0.2) cm (or 13 1 AU), indicating the location where thermal radiation originates from within the dense CSM (Chevalier & Irwin, 2011). This radius, derived from blackbody fits to UV-Optical-NIR data, is typically smaller (as it generally forms deeper) than the radius of photospheric emission () (Moriya et al., 2011) and the surface of last scattering of the dense CSM.
@@ 8.2 Progenitor of SN 2023ixf
Numerous works on SN 2023ixf have estimated the progenitor mass, mass-loss rate, and extent of the CSM encompassing the progenitor. Pre-explosion imaging at the site of SN 2023ixf through HST and Spitzer revealed a point source consistent with an RSG enshrouded by a large amount of dust (Soraisam et al., 2023; Jencson et al., 2023; Neustadt et al., 2024). However, no counterpart was discovered in UV or X-rays (Basu et al., 2023; Matsunaga et al., 2023; Kong, 2023; Panjkov et al., 2023). There is, however, a disparity in the estimates of progenitor mass from the pre-explosion imaging revealing estimates in 2 broad ranges, i.e. (9 – 14 , Kilpatrick et al., 2023; Pledger & Shara, 2023; Van Dyk et al., 2023; Neustadt et al., 2024, and (17 – 22 , Jencson et al., 2023; Soraisam et al., 2023; Niu et al., 2023; Qin et al., 2023. Our numerical hydrodynamical modeling best matched a ZAMS progenitor mass of 10 for SN 2023ixf having a radius of 470 . Only other work that performed hydrodynamical modeling of the complete light curve until the nebular phase also indicated a low-mass progenitor (i.e., 12 , Bersten et al., 2024). The short-plateau nature of SN 2023ixf indicates a relatively lower ejecta mass, which is also reflected in its steep plateau decline rate. Furthermore, the considerable blueshifted offset observed in H ( 3,000 ) during the early phase reinforces this deduction, indicating an escalated degree of stripping undergone by the 10 progenitor of SN 2023ixf.
Pre-explosion observations of the progenitor of SN 2023ixf revealed variability in the mid-IR and near-IR observations from Spitzer and ground-based telescopes (Kilpatrick et al., 2023; Soraisam et al., 2023; Jencson et al., 2023). However, despite this variability, there is no indication of pre-SN outbursts in the pre-explosion imaging conducted by Jencson et al. (2023), nor any signs of variability in the optical spectrum (Dong et al., 2023; Neustadt et al., 2024), which likely denies the existence of episodic mass loss in SN 2023ixf. This lack of pre-explosion outbursts, coupled with the presence of an asymmetric confined CSM (see section 8.4), suggests that the progenitor of SN 2023ixf likely had a rapid rotation and/or underwent a pre-SN interaction with a binary companion (Smith, 2014; Matsuoka & Sawada, 2023). Such an interaction would have led to an enhanced mass loss in the lead-up to the explosion and drive a significant asymmetry in the observed CSM, as inferred from our polarization observations. SN 2023ixf is thus a low-mass RSG progenitor showcasing a multi-faceted CSM geometry arising from enhanced mass loss during its twilight years.
@@ 8.3 Characteristics of circumstellar material around SN 2023ixf
SN 2023ixf showed several signs of an interaction with CSM both photometrically and spectroscopically. The hydrodynamical modeling in Section 7 emphasized the presence of a confined dense CSM responsible for the origin of the flash-ionization features, steep rise in UV flux, bolometric luminosity and temperature, and an extended low-density CSM, which led to the late-phase UV excess and the clumpy features around H. This brings forward the argument that the progenitor of SN 2023ixf had an enhanced wind that developed before the explosion, leading to the delayed shock breakout.
The modeling also revealed that the true extent of the confined CSM is roughly 5 cm (33 AU), more significant than the nearly-flat radius of thermal emission during the early SN evolution in Section 4.3. Our estimates align well with the estimates from comparison with CMFGEN models (Jacobson-Galan et al., 2023), high-cadence early spectroscopy (Bostroem et al., 2023), early light curve modeling (Hiramatsu et al., 2023) and pre-discovery photometry close to the explosion (Li et al., 2023). The confined CSM is characterized by a wind-like structure with a mass-loss rate of 10-2 which is consistent with estimates derived using comparison with CMFGEN models (Jacobson-Galan et al., 2023) and narrow emission features (Zimmerman et al., 2023). The mass-loss rate indicates a progenitor star in the eruptive phase, which could lie anywhere between 10-2 – 0.1 (Smith, 2017). The absence of SN 2023ixf detections in mm-wavelengths by Berger et al. (2023) within the first 4 days, which suggests a mass-loss rate of ,10-2 , aligns with our modeling estimates and is likely influenced by the weak dominance of free-free absorption in the dense CSM along our line-of-sight.
@@
8.4.1 Early polarization evolution: Peak A
Our initial polarization observation (Peak A; 1.1%) at 1.4 d post-explosion aligns with the appearance of flash-ionization lines in the spectral sequence of SN 2023ixf, a consequence of its interaction with the confined dense CSM. Given that the interaction with the dense CSM predominantly influences the luminosity during this phase, the inferred polarization during this period reflects the aspherical geometry of the dense CSM at a position angle of 150 ∘. Early epoch polarization values are consistent with the findings of Maund et al. (2023) and Vasylyev et al. (2023). Since the shock breakout doesn’t happen until 2.2 d, the shock-powered SN ejecta lies within the compact dense CSM at 1.4 d and indicates an asymmetric and confined dense CSM surrounding the progenitor of SN 2023ixf. During this phase, the narrow emission of , arising from the pre-shock CSM, shows a blueshifted and asymmetric feature (Smith et al., 2023), potentially caused by the plane of the dense CSM aligning across our line of sight, obscuring the redshifted portion of the CSM.
Owing to the asphericity of the confined dense CSM and the arguments above, we consider the similarity of the CSM around SN 2023ixf to PTF11iqb (Smith et al., 2015) and SN 1998S (Gerardy et al., 2000), which displayed an equatorial geometry of their CSM (i.e. a disk/torus). Similar to PTF11iqb and SN 1998S, SN 2023ixf exhibited a bright UV peak (see Section 4.2), flash-ionization features (see Section 5.1) and show Peak A in polarization (PA 150∘, see Section 6) within the first four days. However, this phase is relatively short-lived in SN 2023ixf in comparison to PTF11iqb and SN 1998S since the SN ejecta likely envelopes the compact CSM, likely due to its smaller radial extent i.e., ( 30 AU) (see Section 7).

### v2

@@ 7 Light Curve modeling
To estimate the explosion parameters and infer the properties of CSM around SN 2023ixf, we performed numerical light curve modeling of SN 2023ixf. We used one-dimensional multi-frequency radiation hydrodynamics code STELLA (Blinnikov et al., 1998, 2000, 2006) for this purpose. Moriya et al. (2023) computed light-curve models of SNe II with various progenitor masses, explosion energies, and CSM density profile and its extent using STELLA. The model grid adopts the solar-metallicity RSG progenitor models with the ZAMS masses of 10, 12, 14, 16, 18 from Sukhbold et al. (2016) and covers explosion energies from to . The CSM is constructed by assuming the mass-loss rates of with a terminal wind velocity of . The density structure of the wind above the progenitors is characterized by the parameter ranging from to . shapes the velocity/density profile of the wind and is inversely proportional to wind acceleration (Moriya et al., 2017). We refer the reader to Moriya et al. (2023) for additional details regarding the model grid. The extensive details regarding the model comparison will be presented in Moriya & Singh (2024).
@@
Using the pre-computed model grid, we first looked for the best matching models in the g and r bands by searching for models with the minimum . We found that the models from a low-mass progenitor () best fit the light curves of SN 2023ixf. The best fitting models have the explosion energy of around , the mass of around , the mass-loss rate of around 10-2 , the CSM radius of around , and of around . We performed additional numerical light curve calculations to fit the early UV luminosity and found that the model with the CSM radius of matches the observed dataset, including the photospheric velocity (shown in dashed lines in Fig. 13). The mass of the confined CSM below in this model is . This extra model assumes that is uniformly mixed in the entire ejecta. We found that the fully mixed model better fits the transition from the plateau phase to the tail phase. However, they also lead to an excess emission in the redder bands during the late plateau phase as starts diffusing out earlier than in the case of a centrally concentrated 56Ni. A much more prominent effect of -mixing was seen in the case of SN 2009ib (Takáts et al., 2015) and SN 2016gfy (Singh et al., 2019). Since the 56Ni synthesized in SN 2023ixf is high in comparison to normal SNe II (0.03 , Anderson et al., 2014a), the late-plateau luminosity bump is much more prominent in the redder bands of our models.
@@
We observed a notable decline in the UV flux around 40 days in the best-fitting model (having only an inner dense CSM) presented above in Fig.13. To address this decrease and account for the sustained UV excess in the late phase of SN 2023ixf, we modeled the light curves with an extended CSM component situated right outside the inner CSM from to . This additional CSM component possesses a mass of 0.025 and has a density structure proportional to . The low-density extended CSM component helps sustain a prolonged UV brightness until the early nebular phase. Assuming a constant wind velocity of , the outer CSM component is shaped by an average mass-loss rate of 10-4 . The mass-loss rate should increase towards the time of the explosion to form a density structure proportional to . We found that the shallower CSM structure, such as that proportional to , gives a UV flux excess much higher than observed. Using the observed wind velocities of 50 – 115 (Jacobson-Galan et al., 2023; Smith et al., 2023; Zhang et al., 2023) instead of an assumed value, we infer a mass-loss rate of for the inner dense CSM and for the extended CSM.
@@
SN 2023ixf has a higher peak luminosity, higher 56Ni-mass, and a higher photospheric velocity than a typical Type II SN. The peak -band luminosity of SN 2023ixf is -18.2 mag, which is 4 times brighter than compared to the luminosity of a typical Type II SN, i.e., -16.7 mag (Anderson et al., 2014a). The 56Ni-mass of SN 2023ixf is 0.059 0.001 which is 80% higher than the 56Ni-mass estimate of a typical Type II SN, i.e., 0.033 from (Anderson, 2019). Previous studies on SNe II (Hamuy, 2003; Pejcha & Prieto, 2015; Gutiérrez et al., 2017b) have shown that more energetic explosions lead to higher photospheric velocities and a higher 56Ni-yield in the SN. The estimated 56Ni-mass does indicate that SN 2023ixf is a highly energetic event and is in concordance with the high explosion energy of 2 10 estimated from the light curve modeling. Other works on numerical modeling of SN 2023ixf also suggest high explosion energies. Bersten et al. (2024) estimate 1.2 foe as explosion energy for a 12 RSG with 10.9 as the final progenitor mass. Hiramatsu et al. (2023) estimated 1 foe as explosion energy while exploring 2 foe as well for a 12 RSG with 11 as the final progenitor mass.
@@
In our case, where we have the best match with a lower mass progenitor, slightly higher explosion energies must be required to match the observed light curves. Although we fit both photospheric velocities and light curves simultaneously, the mass and radius of the progenitors are fixed since we adopt progenitor models from Sukhbold et al. (2016). Given the inherent degeneracies between the ejecta mass, radius, and explosion energy (Goldberg et al., 2019), our inferred properties likely represent one possible set of solutions rather than a unique one. However, the high-explosion energy cannot entirely explain the peak luminosity of SN 2023ixf. It is enhanced further by early interaction with the confined CSM. The photospheric velocity of SN 2023ixf is 20% faster than a proto-typical Type II SN at 50 d (see Section 5.4), and a power-law fit to photospheric velocity evolution returned exponent -0.47 0.04, which is slower than the average value derived from a large sample of SNe IIP (–0.581 0.034) in Faran et al. (2014). However, since SN 2023ixf is a short-plateau SN, the slightly low-ejecta mass could be one reason for its higher photospheric velocity (Teja et al., 2022).
@@ 8.1 Timeline of significant epochs for all historic observations
Epoch of Polarization Point A | 1.4 d | 4, This paper | |
@@
Peak in UV Luminosity (blue peak) | 4.5 0.5 d | 3, This paper | |
@@
Emergence of Broad absorption | 7 d | 1, 2, 5, 8 | |
End of Flash-Ionisation Phase | 8 d | 1, 5, 8 | |
@@
Epoch of Radio Detection (10 GHz) | 29.2 d | 12 |
(1) Teja et al. (2023a) (2) Smith et al. (2023); (3) Zimmerman et al. (2024); (4) Vasylyev et al. (2023); (5) Jacobson-Galan et al. (2023); (6) Grefenstette et al. (2023); (7) Itagaki (2023); (8) Bostroem et al. (2023); (9) Chandra et al. (2023); (10) Panjkov et al. (2023); (11) Berger et al. (2023); (12) Matthews et al. (2023)
In its infancy, SN 2023ixf showed a rapid evolution spearheaded by the appearance of an increase in ionization in the early flash-ionization phase, accompanied by a rapid ascent in the early UV flux. Additionally, the increase in the color temperature to approximately 35,000 K until around 2.2 days (see Figure 4) signifies that the shock breakout was delayed and occurred within a confined, dense CSM. Our estimate for the epoch of shock breakout differs slightly from that of (Zimmerman et al., 2024). This variation arises because they use the blackbody radius and the shock velocity approximated from the photospheric velocity to compute the extent of the dense CSM. The epoch of peak UV luminosity (and the bolometric peak) at 4.5 d is synonymous with the peak of the He II line flux of Zimmerman et al. (2024), which traces the strength/flux of the flash-ionization features. The change in the line profile of the narrow feature and the drop in its strength seen in the high-resolution spectroscopy of SN 2023ixf (Smith et al., 2023) also happens around 4.4 d, and is synonymous with UV peak. This indicates that the thermal heating and ionization continued beyond the shock breakout (2.2 d). The emergence of a CDS (Chugai, 2009) within the post-shock CSM and the decelerated SN ejecta is probably contributing to the photo-ionization and prolonged heating observed in the flash-ionized features of SN 2023ixf. During the breakout phase, the radius of thermal emission remained relatively constant at (2.0 0.2) cm (or 13 1 AU), indicating the location where thermal radiation originates from within the dense CSM (Chevalier & Irwin, 2011). This radius, derived from blackbody fits to UV-Optical-NIR data, is typically smaller (as it generally forms deeper) than the radius of photospheric emission () (Moriya et al., 2011) and the surface of last scattering of the dense CSM.
@@ 8.2 Progenitor of SN 2023ixf
Numerous works on SN 2023ixf have estimated the progenitor mass, mass-loss rate, and extent of the CSM encompassing the progenitor. Pre-explosion imaging at the site of SN 2023ixf through HST and Spitzer revealed a point source consistent with an RSG enshrouded by a large amount of dust (Soraisam et al., 2023; Jencson et al., 2023; Neustadt et al., 2024). However, no counterpart was discovered in UV or X-rays (Basu et al., 2023; Matsunaga et al., 2023; Kong, 2023; Panjkov et al., 2023). There is, however, a disparity in the estimates of progenitor mass from the pre-explosion imaging revealing estimates in 2 broad ranges, i.e. (9 – 14 , Kilpatrick et al., 2023; Pledger & Shara, 2023; Van Dyk et al., 2023; Neustadt et al., 2024, and (17 – 22 , Jencson et al., 2023; Soraisam et al., 2023; Niu et al., 2023; Qin et al., 2023. Our numerical hydrodynamical modeling best matched a ZAMS progenitor mass of 10 for SN 2023ixf having a radius of 470 . Only other work that performed hydrodynamical modeling of the complete light curve until the early nebular phase indicated a 12 progenitor (Bersten et al., 2024). The short-plateau nature of SN 2023ixf indicates a relatively lower ejecta mass, which is also reflected in its steep plateau decline rate. Furthermore, the considerable blueshifted offset observed in H ( 3,000 ) during the early phase reinforces this deduction, indicating an escalated degree of stripping undergone by the 10 progenitor of SN 2023ixf. A summary of the various estimates of the progenitor mass and the extent and the mass-loss rate of the dense CSM is shown in Figure 15.
Pre-explosion observations of the progenitor of SN 2023ixf revealed variability in the mid-IR and near-IR observations from Spitzer and ground-based telescopes (Kilpatrick et al., 2023; Soraisam et al., 2023; Jencson et al., 2023). However, despite this variability, there is no indication of pre-SN outbursts in the pre-explosion imaging Jencson et al. (2023); Hiramatsu et al. (2023); Ransome et al. (2024), nor any signs of variability in the optical spectrum (Dong et al., 2023; Ransome et al., 2024; Neustadt et al., 2024), which likely denies the existence of episodic mass loss in SN 2023ixf. This lack of pre-explosion outbursts, coupled with the presence of an asymmetric confined CSM (see section 8.4), suggests that the progenitor of SN 2023ixf likely had a rapid rotation and/or underwent a pre-SN interaction with a binary companion (Smith, 2014; Matsuoka & Sawada, 2023). Such an interaction would have led to an enhanced mass loss in the lead-up to the explosion and drive a significant asymmetry in the observed CSM, as inferred from our polarization observations. SN 2023ixf is thus a low-mass RSG progenitor showcasing a multi-faceted CSM geometry arising from enhanced mass loss during its twilight years. The presence of the two-zone CSM in SN 2023ixf (see Section 13) supports the notion of an increased mass-loss rate of the RSG progenitor of SN 2023ixf as it evolved towards its explosion.
@@ 8.3 Characteristics of circumstellar material around SN 2023ixf
SN 2023ixf showed several signs of an interaction with CSM both photometrically and spectroscopically. The hydrodynamical modeling in Section 7 emphasized the presence of a confined dense CSM responsible for the origin of the flash-ionization features, steep rise in UV flux, bolometric luminosity and temperature, and an extended low-density CSM, which led to the late-phase UV excess and the clumpy features around H. This brings forward the argument that the progenitor of SN 2023ixf had an enhanced wind that developed shortly before the explosion, leading to the delayed shock breakout.
The modeling also revealed that the true extent of the confined CSM is roughly 5 cm (33 AU), more significant than the nearly-flat radius of thermal emission during the early SN evolution in Section 4.3. Our estimates align well with the estimates from comparison with CMFGEN models (Jacobson-Galan et al., 2023), high-cadence early spectroscopy (Bostroem et al., 2023), early light curve modeling (Hiramatsu et al., 2023) and pre-discovery photometry close to the explosion (Li et al., 2024). Assuming a wind-velocity of 10 , the confined CSM is characterized by a wind-like structure with a mass-loss rate of 10-2 which is consistent with estimates derived using comparison with CMFGEN models (Jacobson-Galan et al., 2023) and narrow emission features (Zimmerman et al., 2024). The mass-loss rate indicates a progenitor star in the eruptive phase, which could lie anywhere between 10-2 – 0.1 (Smith, 2017). The absence of SN 2023ixf detections in mm-wavelengths by Berger et al. (2023) within the first 4 days, suggests a mass-loss rate of 10-2 , aligns with our modeling estimates and is likely influenced by a strong dominance of free-free absorption in the dense CSM along our line-of-sight.
@@
8.4.1 Early polarization evolution: Point A
Our initial polarization observation (Point A; 1.1%) at 1.4 d post-explosion aligns with the appearance of flash-ionization lines in the spectral sequence of SN 2023ixf, a consequence of its interaction with the confined dense CSM. Given that the interaction with the dense CSM predominantly influences the luminosity during this phase, the inferred polarization during this period reflects the aspherical geometry of the dense CSM at a PA of 150 ∘. Early epoch polarization values are consistent with the findings of Maund et al. (2023) and Vasylyev et al. (2023). Since the shock breakout doesn’t happen until 2.2 d, the shock-powered SN ejecta lies within the compact dense CSM at 1.4 d and indicates an asymmetric and confined dense CSM surrounding the progenitor of SN 2023ixf. During this phase, the narrow emission of , arising from the pre-shock CSM, shows a blueshifted and asymmetric feature (Smith et al., 2023), potentially caused by the plane of the dense CSM aligning across our line of sight, obscuring the redshifted portion of the CSM.
Owing to the asphericity of the confined dense CSM and the arguments above, we consider the similarity of the CSM around SN 2023ixf to PTF11iqb (Smith et al., 2015) and SN 1998S (Gerardy et al., 2000), which displayed an equatorial geometry of their CSM (i.e. a disk/torus). Similar to PTF11iqb and SN 1998S, SN 2023ixf exhibited a bright UV peak (see Section 4.2), flash-ionization features (see Section 5.1) and show Point A in polarization (PA 150∘, see Section 6) within the first four days. However, this phase in SN 2023ixf in similar in comparison to SN 1998S ( 8 d) and longer than that of PTF11iqb ( 3 d) (Jacobson-Galán et al., 2024). This likely indicates the confined CSM in SN 2023ixf has a large radial extent, i.e., 30 AU, Section 7). Assuming an oblate geometry for the CSM, the polarization at Point A corresponds to a major-to-minor axis ratio of 1.2 (Hoflich, 1991).